En. Iman is the owner of 100 bonds issued by LYC Bhd. These bonds have 8 years remaining to maturity, an annual coupon payment of RM80, and a par value of RM1,000. He purchased the bond 2 years ago when the interest rate was 8.5 percent. Due to the pandemic, the interest rate has fallen by 1.7 percent from it was 2 years ago. Considering this, En Iman decided not to hold the bonds till maturity but to sell it today.



Help him to compute his total capital gain from this investment

1
Expert's answer
2022-01-04T10:20:48-0500

price of a bond:

"P=\frac{PMT(T_1)}{1+r}+\frac{PMT(T_2)}{(1+r)^2}+...+\frac{PMT(T_n)+FV}{(1+r)^n}"


where PMT(tn) is Coupon Payment at Time n,

n is number of periods,

r is interest rate,

FV is par value


then we have, for the first two years:

"P_1=\frac{80}{1+0.085}+\frac{80+1000}{(1+0.085)^2}=991.14"


for the last 8 years:

"P_2=\frac{80}{1+0.017}+\frac{80}{(1+0.017)^2}+...+\frac{80+1000+991.14}{(1+0.017)^8}=2325.52"


total capital gain:

"100P_2=232552"